# On inverse problem for generalized Kerr-type nonlinearities

### v1

@@ 3.3. Test data: Complex geometrical optics (CGO) solution
Proof of Lemma 3.4. Substituting the ansatz $v=e^{-\frac{\rho\cdot x}{h}}(1+r_{\rho})$ into the equation $(\Delta+q_{0})v=0$ in $\Omega$, we obtain
00 0 $=h^{2}e^{\frac{\rho\cdot x}{h}}(\Delta+q_{0})\left(e^{-\frac{\rho\cdot x}{h}}(1+r_{\rho})\right)$
@@
$h^{2}e^{\frac{\omega\cdot x}{h}}(\Delta+q_{0})(e^{-\frac{\omega\cdot x}{h}}\tilde{r}_{\boldsymbol{\beta}})=-h^{2}e^{-\mathbf{i}\frac{\omega\cdot x}{h}}\left(\partial^{\boldsymbol{\beta}}q_{0}+\sum_{\boldsymbol{\gamma}\leq\boldsymbol{\beta},\boldsymbol{\gamma}\neq 0}(\partial^{\boldsymbol{\gamma}}q_{0})(\partial^{\boldsymbol{\beta}-\boldsymbol{\gamma}}r_{\rho})\right).$
@@
$\partial_{j}\partial_{k}v_{h}=e^{-\frac{J}{h}}e^{-\frac{1}{h}(\rho\cdot x-t)}\left(\left(\frac{\rho_{j}}{h}\right)\left(\frac{\rho_{k}}{h}\right)(1+r_{\rho})+\left(\frac{\rho_{k}}{h}\right)\partial_{j}r_{\rho}+\left(\frac{\rho_{j}}{h}\right)\partial_{k}r_{\rho}+\partial_{j}\partial_{k}r_{\rho}\right),$
@@
then for all sufficiently small $h\in(0,1)$, we have $f_{h}\in U_{\delta_{0}}$. This ensures the CGO solution Lemma 3.5 can be substituted into the indicator functionals $\mathcal{I}$ and $\widetilde{\mathcal{I}}$.
@@ 3.4. Conclusion
Theorem 3.6 (see also Theorem 1.1 above).
Let $\Omega\subset\mathbb{R}^{n}$ ($n\geq 2$) be a bounded $C^{1,1}$-domain. Assume there exist $0\geq q_{0}\in W^{\lfloor\frac{n}{2}\rfloor+1,\infty}(\Omega)$, $q_{1}\in L^{\infty}(\Omega)$ with (1.5), a constant $C_{*}>0$, an integer $\alpha_{1}\in\mathbb{Z}_{\geq 2}$, a real number $\alpha_{2}>\alpha_{1}$, such that the nonlinear term $q$ satisfies (1.1b). Let $t_{*}:=\inf_{x\in D}\omega\cdot x$ and let $f_{h}$ be the function given in Lemma 3.5. For each $\omega\in\mathbb{S}^{n-1}$, we choose
(3.6) $J>\max\left\{\frac{B(\omega)-b(\omega)(2+2\alpha_{1})}{\alpha_{1}},\frac{(B(\omega)-b(\omega))(2+\alpha_{2})}{\alpha_{2}-\alpha_{1}}\right\}\quad\text{and}\quad b(\omega)<t<B(\omega).$
@@
    $\lim_{h\searrow 0}e^{\frac{(\alpha_{1}+2)J}{h}}\lvert\mathcal{I}(f_{h})\rvert=+\infty.$
@@
Theorem 3.6 implies that for any fixed direction $\omega\in\mathbb{S}^{n-1}$, one can determine whether the hyperplane ${\omega\cdot x=t}$ intersects the inclusion $D$ by testing the limiting behavior of $e^{\frac{(\alpha_{1}+2)J}{h}}\lvert\mathcal{I}(f_{h})\rvert$ as $h\to 0$. By varying $t$, we can identify the critical value $t_{*}$ such that the hyperplane ${\omega\cdot x=t_{*}}$ just touches $D$, although the limiting behavior at $t=t_{*}$ itself is not known. By applying this procedure in every direction $\omega\in\mathbb{S}^{n-1}$, we reconstruct the convex hull of $D$ and thereby conclude Theorem 1.1.
Let us first make some observations before proving Theorem 3.6. Using Lemma 3.3, we insert the CGO solution $v_{h}$ given in (3.5) into $\widetilde{\mathcal{I}}$ to see that
~$\widetilde{\mathcal{I}}(f_{h})=-\int_{D}q_{1,D}(x)\lvert v_{h}\rvert^{2+\alpha_{1}}\,\mathrm{d}x-\int_{\Omega}R(x,\lvert v_{h}\rvert)\lvert v_{h}\rvert^{2}\,\mathrm{d}x,$
@@
$\quad+e^{\frac{(\alpha_{1}+2)J}{h}}(\mathcal{I}(f_{h})-\widetilde{\mathcal{I}}(f_{h}))$
By using Lemma 3.2, Lemma 3.5 and (3.6), we see that
@@
We now ready to prove our main result Theorem 3.6.
Proof of Theorem 3.6(i). Using (3.4) we estimate
@@
Proof of Theorem 3.6(ii). Using (1.5b), we obtain
@@
$e^{\frac{(\alpha_{1}+2)J}{h}}\left|\int_{D}q_{1,D}(x)\lvert v_{h}\rvert^{2+\alpha_{1}}\,\mathrm{d}x\right|\geq 2^{-2-\alpha_{1}}\mu\int_{D}e^{-\frac{J}{h}(2+\alpha_{1})}e^{-\frac{2+\alpha_{1}}{h}(\omega\cdot x-t)}\,\mathrm{d}x.$

### v2

@@ 3.3. Test data: Complex geometrical optics (CGO) solution
Proof of Lemma˜3.4. Substituting the ansatz $v=e^{-\frac{\rho\cdot x}{h}}(1+r_{\rho})$ into the equation $(\Delta+q_{0})v=0$ in $\Omega$, we obtain
$0$ $=h^{2}e^{\frac{\rho\cdot x}{h}}(\Delta+q_{0})\left(e^{-\frac{\rho\cdot x}{h}}(1+r_{\rho})\right)$
@@
$h^{2}e^{\frac{\omega\cdot x}{h}}(\Delta+q_{0})(e^{-\frac{\omega\cdot x}{h}}\tilde{r}_{\boldsymbol{\beta}})=-h^{2}e^{-\mathbf{i}\frac{\omega^{\perp}\cdot x}{h}}\left(\partial^{\boldsymbol{\beta}}q_{0}+\sum_{\boldsymbol{\gamma}\leq\boldsymbol{\beta},\boldsymbol{\gamma}\neq 0}(\partial^{\boldsymbol{\gamma}}q_{0})(\partial^{\boldsymbol{\beta}-\boldsymbol{\gamma}}r_{\rho})\right).$
@@
$\partial_{j}\partial_{k}v_{h}=e^{-\frac{J}{h}}e^{-\frac{1}{h}(\rho\cdot x-t)}\left(\left(\frac{\rho_{j}}{h}\right)\left(\frac{\rho_{k}}{h}\right)(1+r_{\rho})-\left(\frac{\rho_{k}}{h}\right)\partial_{j}r_{\rho}-\left(\frac{\rho_{j}}{h}\right)\partial_{k}r_{\rho}+\partial_{j}\partial_{k}r_{\rho}\right),$
@@
then for all sufficiently small $h\in(0,1)$, we have $f_{h}\in U_{\delta_{0}}$. This ensures the CGO solution Lemma˜3.5 can be substituted into the indicator functionals $\mathcal{I}$ and $\widetilde{\mathcal{I}}$.
@@ 3.4. Conclusion
Theorem 3.6 (see also Theorem˜1.1 above).
Let $\Omega\subset\mathbb{R}^{n}$ ($n\geq 2$) be a bounded $C^{1,1}$-domain. Assume there exist $0\geq q_{0}\in W^{\lfloor\frac{n}{2}\rfloor+1,\infty}(\Omega)$, $q_{1}\in L^{\infty}(\Omega)$ with (1.5), a constant $C_{*}>0$, an integer $\alpha_{1}\in\mathbb{Z}_{\geq 2}$, a real number $\alpha_{2}>\alpha_{1}$, such that the nonlinear term $q$ satisfies (1.1b). Let $t_{*}:=\inf_{x\in D}\omega\cdot x$ and let $f_{h}$ be the function given in Lemma˜3.5. For each $\omega\in\mathbb{S}^{n-1}$, we choose
(3.6) $J>\max\left\{\frac{(B(\omega)-b(\omega))(2+2\alpha_{1})}{\alpha_{1}},\frac{(B(\omega)-b(\omega))(2+\alpha_{2})}{\alpha_{2}-\alpha_{1}}\right\},\quad b(\omega)<t<B(\omega).$
@@
    $\lim_{h\searrow 0}e^{\frac{(\alpha_{1}+2)J}{h}}\lvert\mathcal{I}(f_{h})\rvert$$\infty$.
@@
Theorem˜3.6 implies that for any fixed direction $\omega\in\mathbb{S}^{n-1}$, one can determine whether the hyperplane $\omega\cdot x=t$} intersects the inclusion $D$ by testing the limiting behavior of $e^{\frac{(\alpha_{1}+2)J}{h}}\lvert\mathcal{I}(f_{h})\rvert$ as $h\to 0$. By varying $t$, we can identify the critical value $t_{*}$ such that the hyperplane $\omega\cdot x=t_{*}$} just touches $D$, although the limiting behavior at $t=t_{*}$ itself is not known. By applying this procedure in every direction $\omega\in\mathbb{S}^{n-1}$, we reconstruct the convex hull of $D$ and thereby conclude Theorem˜1.1.
Let us first make some observations before proving Theorem˜3.6. Using Lemma˜3.3, we insert the CGO solution $v_{h}$ given in Lemma˜3.5 into $\widetilde{\mathcal{I}}$ to see that
$\widetilde{\mathcal{I}}(f_{h})=-\int_{D}q_{1,D}(x)\lvert v_{h}\rvert^{2+\alpha_{1}}\,\mathrm{d}x-\int_{\Omega}R(x,\lvert v_{h}\rvert)\lvert v_{h}\rvert^{2}\,\mathrm{d}x.$
@@
$\quad+e^{\frac{(\alpha_{1}+2)J}{h}}(\mathcal{I}(f_{h})-\widetilde{\mathcal{I}}(f_{h})).$
By using Lemma˜3.2, Lemma˜3.5 and (3.6), we see that
@@
We now ready to prove our main result Theorem˜3.6.
Proof of Theorem˜3.6(i). Using Lemma˜3.5 and (3.4) we estimate
@@
Proof of Theorem˜3.6(ii). Using (1.5b), we obtain
@@
$e^{\frac{(\alpha_{1}+2)J}{h}}\left|\int_{D}q_{1,D}(x)\lvert v_{h}\rvert^{2+\alpha_{1}}\,\mathrm{d}x\right|\geq 2^{-2-\alpha_{1}}\mu\int_{D}e^{-\frac{2+\alpha_{1}}{h}(\omega\cdot x-t)}\,\mathrm{d}x.$